given: f(a)=(sqrt)a^2+3a, find a(-2)

Thanks!

Answer by stanbon(75887)

f(a)=(sqrt)a^2+3a, find a(-2)
f(-2)=sqrt((-2)^2)+3(-2)
f(-2)= sqrt(4)-6
f(-2)=2-6
f(-2)=-4
Cheers,
Stan H.